OpenStudy (ttop0816):

Perform the indicated operation. 2 3/8 · 5 3/4 19/46 13 19/46 13 21/32

OpenStudy (cwrw238):

you can do this by first converting each mixed number to an improper fraction 2 3/8:- 2 = 16/8 16/8 + 3/8 = 19/8 can you perform a similar operation on 5 3/4 ?

OpenStudy (mathstudent55):